A dockworker applies a constant horizontal force o
Answer #1

Given F = ma = 70 N

displacement s = 15 m

time t = 6 s

initial speed u = 0m/s

from s = ut + 0.5*a*t^2

15 = 0.5 * a * 36

a = 0.83 m/s^2

a) there fore m = F/a = 70/0.83

mass m = 84.33 kg

b) After the removel of applied force the block
moves with constant velocity

after 6 s the velocity is

v = u + at

v = 0 + 0.83 * 6

v = 4.98 m/s

distance d = v* t = 4.98 * 6

v = 29.88 m
Answer #2

SOLUTION ;


Ice block moves 15 m in 6 sec. Initial speed is zero.


Hence, using formula :

s = ut + 1/2 a t^2 

=> 15 = 0 + 1/2 a (6)^2

=> 15 = 18 a 

=> a = 15/18 = 5/6 m/sec^2 


a.


So,


Mass of the block, m 

= F/a 

= 70 / (5/6) kg 

= 84 kg 


Mass of ice block is = 84 kg (ANSWER).


b.


Speed of block at t = 6 sec :

= v6 + at = 0 + 5/6 * 6 = 5 m/sec.


So distance moved in next 6 sec ;

= v6 * (t) + 1/2 (a) (t)^2

= 5 (6) + 1/2 (5/6) (6)^2

= 30 + 15

= 45 m 


Distance moved in next 6 sec = 45 m (ANSWER)







answered by: Tulsiram Garg

> In part (b) , 2nd line, please read " = v6 + at " as " v6 = u + at "

Tulsiram Garg Sun, Oct 3, 2021 6:55 AM
Answer #3

I am correcting part (b) as under :


b.


Speed of block at t = 6 sec :

v6 = u + at = 0 + (5/6) * 6 = 5 m/s



So distance moved in next 6 sec when force is no more applied : 

= v6  * (t) 

= 5 * 6

= 30 m 


Distance moved in next 6 sec = 30 m (ANSWER)



answered by: Tulsiram Garg